The most recent financial statements for Bello Co. are shown here: Income Statement Sales Costs Taxable Balance Sheet Curent 12,040 Debt 16,720 14,200 Fixed assets 33,750 Eqity 29,070 assets $6,400 $45,790 Total $45,790 income Taxes (23%) 1,472 Net income $ 4,928 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 35 percent dividend payout ratio. What is the internal growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded 2 decimal places, e.g., 32.16.) Internal growth rate
Answer #1

ROA=Net income/Total assets

=(4928/45790)=0.107621751

Retention ratio=1-payout ratio

=(1-0.35)=0.65

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.107621751*0.65)/[1-(0.107621751*0.65)]

=7.52%(Approx).